# A Survey on Integration of Large Language Models with Intelligent Robots

### v2

@@ 1 Introduction
Over the last decade, we have witnessed remarkable progress in the field of robotics in applying language models (LMs). This progress includes not only human-like communication but also understanding and reasoning capabilities of robots thereby significantly improving their effectiveness across various tasks, from household chores to industrial operations [105, 52]. In the early stage of work, the success stemmed from statistical model analyzing and predicting words in linguistic expressions. These models enable robots to interpret human commands [110, 121], understand contexts [5, 3], represent the world [50], and interact with humans [134], albeit with a limited depth of understanding. Then, the adoption of Transformer architecture with self-attention mechanisms [140], particularly pre-trained LMs such as BERT [27], has elevated the capability of capturing complex patterns while fine-tuning models for specific tasks. However, the performance of these models is often contingent upon limited datasets, constraining their ability to grasp deeper contextual understanding and generalize across diverse scenarios.
@@
With the advancement of large language models (LLMs), language-based robotics introduce innovative changes across various domains such as information retrieval, reasoning tasks, adaptation to environments, continuous learning and improvements, etc [64, 61]. These LLMs, characterized by their vast parameter sizes and training on internet-scale datasets, offer zero- and few-shot learning capabilities for downstream tasks without requiring additional parameter updates. These prominent advancements come from emergent abilities, defined as “the abilities that are not present in small models but arise in large models” in the literature [147]. The abilities have significantly enhanced robots’ performance in understanding, inferring, and responding to open-set instructions by leveraging extensive common-sense knowledge [9]. Furthermore, prompt creation technologies, called prompt engineering, have enabled LLMs to incorporate richer contextual information through free-form language descriptions or interactive dialogues, facilitating generalized reasoning [148]. The introduction of in-context learning abilities [9] leads LLMs to generate outputs in expected formats, such as JSON, YAML, or PDDL, or even code, based on provided instructions or demonstrations in prompts [42, 87]. Recent LLMs, such as GPT-4, have further expanded capabilities by integrating with external robotics tools such as planners or translators [90].
@@
Despite the diverse capabilities of LLMs, their utilization faces several challenges [69]. Firstly, LLMs often generate inaccurate or unexpected responses. As the safety of robot execution is the most important deployment factor, LLM-based robotic applications require filtering and correction mechanisms to ensure safety. Second, the emergent abilities, such as in-context learning, are not predictable and consistent yet [20]. Even minor alterations to input text may lead to unpredictable changes in response. Third, a well-designed prompt enables robots to effectively leverage the abilities of LLMs but there is a lack of systematic guidelines supporting key components of robotic systems, hindering seamless integration [36, 164, 54]. Therefore, we need to investigate component-wise LLM engagements in robotics toward an understanding of limitations and safety.
@@
In addition, we provide comprehensive guidelines and examples for prompt engineering, aimed at enabling beginners to access LLM-based robotics solutions. Our tutorial-level examples illustrate how fundamental functionalities of robotic components can be augmented or replaced by introducing four types of exemplary prompts: conversational prompt for interactive grounding, directive prompt for scene-graph generation, planning prompt for few-shot planning, and code-generation prompt for reward generation. By providing rules and tips for prompt construction, we outline the process of generating well-designed prompts to yield outputs in the desired format in the desired format. These principles ensure effective LLM-guided enhancements in robotics applications, without parameter adjustments.
@@
Behavior, tier=level3, fill=brown!20, for tree=fill=fluorescentpink!20, draw=fluorescentpink, child2 [Passive, tier=level4, fill=brown!20, for tree=fill=fluorescentpink!20, draw=fluorescentpink, child3 [ PG-InstructBLIP [37], TidyBot [151], ViLa [53], LAN-grasp [104], RT-Grasp [155], ConceptGraphs [41], VoxPoser [59], LM-Nav [122], FM-Loc [103], ConceptFusion [66], LLM-Grounder [157], WALL-E [143] , tier=level5, fill=fluorescentpink!0, draw=fluorescentpink, child4 ] ] [Active, tier=level4, fill=brown!20, for tree=fill=fluorescentpink!20, draw=fluorescentpink, child3 [ LLM-Planner [129], Kwon et al. [79], Matcha [168], MultiPLY [48] ORION [26] , tier=level5, fill=fluorescentpink!0, draw=fluorescentpink, child4 ] ] ] ] [Planning, tier=level2, for tree=fill=fluorescentpurple!20, draw=fluorescentpurple, child1 [Task
@@
Planning, tier=level4, fill=brown!20, for tree=fill=fluorescentpurple!20, draw=fluorescentpurple, child2 [ Swarm-GPT [68], VoxPoser [59] , tier=level5,fill=fluorescentpurple!0, draw=fluorescentpurple, child4 ] ] [ Kwon et al. [80], RoCo [97], [152], LLM-GROP [30], AutoTAMP [19], Scaling Up and Distilling Down [44], Text2Motion [89] , tier=level5, fill=fluorescentpurple!0, draw=fluorescentpurple, child4 ] ] ] [Control, tier=level2, for tree=fill=fluorescentorange!20, draw=fluorescentorange, child1 [Direct
@@ 2 Preliminary
We briefly review language models in two parts, before and after the advent of LLM. Unlike the overview in the previous literature [164], we limit the period for LM, which is before the advent of LLM, to the period of neural language models when recurrent neural networks (RNNs) [34] started to be used and models such as GPT-2 [115] with Transformers were used. We also provide a brief explanation of LLMs with related terminologies and techniques.
@@ 2.1 Language Models in Robotics
Research in LM-based robotics has primarily explored neural language models, used for sequential data processing. In the early stage, RNN-based LMs [46, 24] transform linguistic commands into a sequence of actions [100, 7] or formal languages [40] by leveraging RNN’s sequence-to-sequence modeling capabilities. Using an RNN as a linguistic encoder, LMs also translate textual inputs into linguistic features, which could then be mapped to visual features for referring object identification [121, 125]. Nonetheless, the long-term dependency problem in RNN restricts their application scope. Then, the introduction of the Transformer architecture [140] overcomes these limitations, enabling its application in tasks requiring long-range comprehension, such as vision-language navigation [17, 15].
Prior approaches heavily depend on training datasets, lowering generalization abilities. However, the scalability of Transformer-based models and self-supervised learning techniques, such as masked language modeling, has led to the development of internet-scale pre-trained models, such as BERT [27] or GPT-2 [115]. These pre-trained models demonstrate general linguistic understanding abilities. This advancement has allowed for fine-tuning these models for specific tasks [74, 75, 124]. Moreover, the utilization of pre-trained multimodal language models, such as CLIP [116], facilitates the exploitation of aligned features across modalities, such as vision and language, enriching the application in robotic studies [76, 126].
@@ 2.2 Large Language Model in Robotics
Recent advancements in LLMs, such as GPT-3 [9], GPT-4 [107], LLaMA [136], Llama 2 [137], and Gemini [2], demonstrate notable improvements in understanding, contextual awareness, generalization capabilities, and knowledge richness, surpassing earlier language models. These improvements are from their training on vast datasets with billion-scale parameters, enabling them to capture intricate data patterns. Additionally, advanced learning strategies, such as reinforcement learning from human feedback, have been developed to align the behaviors of LLMs with human values or preferences [108].
Alternatively, prompt engineering, leveraging in-context learning (ICL) [9], represents another significant advancement, allowing LLMs to directly learn from prompts without additional training. The effectiveness of the prompt engineering hinges on the prompt’s design and quality, with enhancements including detailed task descriptions, few-shot examples, or more model-digestible formats (e.g., ### as a stop symbol [167]). Furthermore, the chain-of-thought (CoT) prompting method [148] emerges, incorporating intermediate reasoning steps into prompts, resulting in significant enhancement on complex reasoning tasks. Ongoing research endeavors to further improve the reasoning abilities of LLMs, with developments such as tree-of-thought [160] or graph-of-thought [6].
On the other hand, various LLM-based robotics studies have tried directly training LLMs [8, 170]. However, full fine-tuning, training the entire model on the task-specific data, is not only computationally expensive but also costly to obtain enough data, due to their large-scale parameters. To address these issues, researchers have developed parameter-efficient fine-tuning methods, such as adapters—small, trainable networks inserted into each layer of an LLM for task-specific tuning [49], and LoRA [51], which applies a low-rank constraint to approximate the updated matrix in each layer.
These developments in LLMs are significantly influencing robotics, setting the stage for a deeper exploration of LLM applications within robotic systems.
@@ 3 Communication
We investigate the utilization of LLMs to facilitate human-like communication in robotics, enabling robots to interact with humans and other robotic agents effectively [98]. We categorize the communication capabilities into two primary areas: (1) language understanding and (2) language generation, as detailed in Fig. 1, which shows the detailed categorization alongside relevant studies, referred in green cells.
@@ 3.1 Language Understanding
We review language understanding capabilities, addressing how LLMs handle the variability and ambiguity of linguistic inputs through interpretation and grounding.
Interpretation involves transforming natural-language inputs into robot-operatable semantic representations, ranging from formal languages, such as linear temporal logic (LTL) [94, 159] and planning domain definition language (PDDL) [154, 90, 19, 42], to programming languages, such as Python [56, 76]. To aid in interpreting free-form sentences, researchers leverage LLMs’ ICL capabilities, providing guidelines and demonstrations within prompts [122, 56, 90, 76]. Despite the efforts, LLMs often fail to satisfy syntax or capture precise semantics when translating an input into formal languages. Solutions include simplifying vocabulary or fine-tuning LLMs with domain-agnostic data [94, 159]. Translation systems, such as Lang2LTL [92], exemplify how LLMs translate landmark-referring expressions in navigational commands into LTL symbols. Further improvements often involve using human feedback and syntax checkers to correct generated formal language translations [19, 42]. For instance, Guan et al. [42] present a human-in-the-loop translation framework, in which human domain experts repeatedly review PDDL descriptions and provide feedback in natural language.
Grounding is to map linguistic expressions to referents recognizable by robots, such as behaviors or objects. Early studies find the mapping by maximizing the cosine similarity between the word embedding of LLM outputs and real-world referents [58, 117, 94, 76]. Subsequent studies incorporate commonsense knowledge from LLMs for contextual support in grounding linguistic label of objects [118, 41]. For instance, Gu et al. [41] demonstrate how LLMs can ground ‘something to use as a paperweight’ to a ceramic vase based on size and weight assumptions we know. However, grounding accuracy depends on the detail and accuracy of the world model. To address this, researchers augment LLMs with multimodal capabilities to directly correlate linguistic inputs with sensory percepts [32, 158, 47, 114], or enable LLMs to interact with environments [157, 168] or humans [61, 120, 109] for better context gathering. For instance, a 3D visual grounding method, LLM-Grounder [157], actively gathers environmental information using vision tools, such as LERF [72] and OpenScene [111].
@@ 3.2 Language Generation
Task-dependent language generation focuses on producing language with specific functional objectives, being declarative or imperative. To generate open-ended declarative statements, researchers often provide LLMs with contextual information [21, 97, 62]. However, LLMs often result in repetitive and factually inconsistent outputs, confined by the reliance on previous dialogues and commonsense knowledge [21, 84]. Consequently, researchers augment LLMs with auxiliary knowledge sources to broaden the scope of available information [4, 156, 22]. For instance, Axelsson and Skantze [4] enhance a robot museum guide with knowledge graphs. Furthermore, researchers instruct LLMs to clarify ambiguities by generating imperative instructions requesting human assistance [61, 26]. To improve inference steps, probabilistic models have been introduced to evaluate the uncertainty of situations [120, 109]. For instance, KnowNo [120] and CLARA [109] interaction systems assess confidence and semantic variance, respectively, triggering generation only when these metrics indicate significant uncertainty.
Task-independent language generation involves crafting expressions with social-emotional objectives [12], by embedding non-verbal cues (e.g., non-verbal sounds, hand gestures, and facial expressions) within prompts to enhance engagement and empathy [73, 81]. For example, Khoo et al. [73] have developed a conversational robot that generates empathetic responses using transcribed audio and visual cues. However, conversations with LLMs remain superficial due to the limited knowledge and dialogue history [65]. To overcome this, researchers integrate memory modules into LLMs, enabling them to distill and store information from conversations in a structured format [65, 63, 23, 162]. For example, the companion robot designed by Irfan et al. [65] continuously updates the robot’s memory based on interactions with users to generate personalized dialogues.
@@ 4.1 Sensing Modalities
Visual perception tasks involve the interpretation of visual information such as image or point clouds. Pre-trained visual-language models (VLMs), such as CLIP [116] and InstructBLIP [83], allows LLM-based robots to directly utilize image sources. For instance, recent LLM-based manipulation systems, such as TidyBot [151] and RoCo [97], use image-inferred object labels or scene descriptions, generated from CLIP and OWL-ViT [101], respectively. In addition, researchers extend reasoning capabilities by applying VLMs on downstream tasks such as image captioning [41] and visual question answering (VQA) [37, 79, 104]. The downstream tasks enable LLMs to subsequently request VLMs to infer object properties (e.g., material, fragility) [37] or ground object parts for grasping [104]. However, images are often challenging to acquire spatial-geometric information.
Alternatively, Huang et al. [56] associate visual-language features from a VLM (i.e., LSeg [82]) with three-dimensional (3D) point clouds for 3D map reconstruction. Further, Jatavallabhula et al. [66] improve this association mechanism with RGB-D images by introducing fine-grained and pixel-aligned features from VLMs. However, association with 3D information tends to be memory intensive, limiting scalability for large scenes [56, 66, 157]. As an alternative solution, researchers often associate geometric and semantic features with 3D scene graphs [41].

### v3

@@ 1 Introduction
Over the last decade, we have witnessed remarkable progress in the field of robotics in applying language models (LMs). This progress includes not only human-like communication but also understanding and reasoning capabilities of robots, thereby significantly improving their effectiveness across various tasks, from household chores to industrial operations [105, 52]. In the early stage of work, the success stemmed from statistical models analyzing and predicting words in linguistic expressions. These models enable robots to interpret human commands [110, 121], understand contexts [5, 3], represent the world [50], and interact with humans [135], albeit with a limited depth of understanding. Then, the adoption of Transformer architecture with self-attention mechanisms [141], particularly pre-trained LMs such as BERT [26], has elevated the capability of capturing complex patterns while fine-tuning models for specific tasks. However, the performance of these models is often contingent upon limited datasets, constraining their ability to grasp deeper contextual understanding and generalize across diverse scenarios.
@@
With the advancement of large language models (LLMs), language-based robotics introduce innovative changes across various domains such as information retrieval, reasoning tasks, adaptation to environments, continuous learning and improvements, etc [64, 61]. These LLMs, characterized by their vast parameter sizes and training on internet-scale datasets, offer zero- and few-shot learning capabilities for downstream tasks without requiring additional parameter updates. These prominent advancements come from emergent abilities, defined as “the abilities that are not present in small models but arise in large models” in the literature [148]. The abilities have significantly enhanced robots’ performance in understanding, inferring, and responding to open-set instructions by leveraging extensive common-sense knowledge [8]. Furthermore, prompt engineering techniques have enabled LLMs to incorporate richer contextual information through free-form language descriptions or interactive dialogues, facilitating generalized reasoning [149]. The introduction of in-context learning abilities [8] leads LLMs to generate outputs in expected formats, such as JSON, YAML, or PDDL, or even code, based on provided instructions or demonstrations in prompts [42, 86]. Recent LLMs, such as GPT-4, have further expanded capabilities by integrating with external robotics tools such as planners or translators [89].
@@
Despite the diverse capabilities of LLMs, their utilization faces several challenges [69]. Firstly, LLMs often generate inaccurate or unexpected responses. As the safety of robot execution is one of the most important deployment factors, LLM-based robotic applications require filtering and correction mechanisms to ensure safety. Second, the emergent abilities, such as in-context learning, are not predictable and consistent yet [19]. Even minor alterations to input text may lead to unpredictable changes in response. Third, a well-designed prompt enables robots to effectively leverage the abilities of LLMs but there is a lack of systematic guidelines supporting key components of robotic systems, hindering seamless integration [35, 164, 54]. Therefore, we need to investigate component-wise LLM engagements in robotics toward an understanding of limitations and safety.
@@
In addition, we provide comprehensive guidelines and examples for prompt engineering, aimed at enabling beginners to access LLM-based robotics solutions. Our tutorial-level examples illustrate how fundamental functionalities of robotic components can be augmented or replaced by introducing four types of exemplary prompts: conversational prompt for interactive grounding, directive prompt for scene-graph generation, planning prompt for few-shot planning, and code-generation prompt for reward generation. By providing rules and tips for prompt construction, we outline the process of generating well-designed prompts to yield outputs in the desired format. These principles ensure effective LLM-guided enhancements in robotics applications, without parameter adjustments.
@@
Behavior, tier=level3, fill=brown!20, for tree=fill=fluorescentpink!20, draw=fluorescentpink, child2 [Passive, tier=level4, fill=brown!20, for tree=fill=fluorescentpink!20, draw=fluorescentpink, child3 [ PG-InstructBLIP [37], TidyBot [152], ViLa [53], LAN-grasp [103], RT-Grasp [156], ConceptGraphs [41], VoxPoser [59], LM-Nav [122], FM-Loc [102], ConceptFusion [66], LLM-Grounder [158], WALL-E [144] , tier=level5, fill=fluorescentpink!0, draw=fluorescentpink, child4 ] ] [Active, tier=level4, fill=brown!20, for tree=fill=fluorescentpink!20, draw=fluorescentpink, child3 [ LLM-Planner [129], Kwon et al. [78], Matcha [168], MultiPLY [48] ORION [25], LLM-POP [132] , tier=level5, fill=fluorescentpink!0, draw=fluorescentpink, child4 ] ] ] ] [Planning, tier=level2, for tree=fill=fluorescentpurple!20, draw=fluorescentpurple, child1 [Task
@@
Planning, tier=level4, fill=brown!20, for tree=fill=fluorescentpurple!20, draw=fluorescentpurple, child2 [ Swarm-GPT [68], VoxPoser [59] , tier=level5,fill=fluorescentpurple!0, draw=fluorescentpurple, child4 ] ] [ Kwon et al. [79], RoCo [96], Xia et al. [153], LLM-GROP [29], AutoTAMP [18], Scaling Up and Distilling Down [44], Text2Motion [88] , tier=level5, fill=fluorescentpurple!0, draw=fluorescentpurple, child4 ] ] ] [Control, tier=level2, for tree=fill=fluorescentorange!20, draw=fluorescentorange, child1 [Direct
@@ 2 Preliminary
We briefly review language models used in robotics, categorizing them in terms of pre- and post-LLM eras. Unlike previous literature [164], we define the pre-LLM era as the period for LMs up to the advent of GPT-2 [115], characterized by neural language models such as recurrent neural networks (RNNs) [33] and early Transformer architectures. We then provide a brief explanation of LLMs, introducing terminologies and techniques for subsequent reviews.
@@ 2.1 Language Models in Robotics
In the pre-LLM era, early-stage studies have primarily focused on sequential data processing, using RNN-based models [46, 23]. The models are often to transform linguistic commands into a sequence of actions [99, 6] or formal languages [40], leveraging RNN’s sequence-to-sequence modeling capabilities. On the other hand, researchers have also used RNNs as language encoders converting textual input into linguistic features that could be mapped to visual features for referring object identification [121, 125]. However, the long-term dependency issue in RNNs restricts the scope of applications. Then, the Transformer architecture [141], which is a non-sequential model supporting long-range comprehension, has enabled new robotic tasks, such as vision-and-language navigation [16, 14].
The studies in the pre-LLM era also show improved application performance unlike previous methods trained on small, task-specific datasets. Transformer-based models and self-supervised learning techniques, such as masked language modeling, have led to the development of internet-scale pre-trained models, including BERT [26] and GPT-2 [115]. These models exhibit a broad understanding of language, enabling both (1) improved generalization abilities and (2) fine-tuning for specific robotic tasks [74, 75, 124]. In addition, researchers have also developed LMs that process multi-modal information [116] since robotic applications often require accessing diverse multimodalities, such as natural language and vision, for interaction with users and environment [76, 126].
@@ 2.2 Large Language Model in Robotics
Recent advancements in LLMs, such as GPT-3 [8], GPT-4 [107], LLaMA [137], Llama 2 [138], and Gemini [2], demonstrate notable improvements in understanding, contextual awareness, generalization capabilities, and knowledge richness, surpassing earlier language models. These improvements are from their training on vast datasets with billion-scale parameters, enabling them to capture intricate data patterns. Further, advanced learning strategies, such as reinforcement learning from human feedback, have been developed to align the behaviors of LLMs with human values or preferences [108]. However, learning with large-scale parameters requires computationally expensive costs of updating the entire model. To address this issue, researchers have developed parameter-efficient fine-tuning methods (e.g., an adapter [49] and LoRA [51]) for robotic tasks. For example, LLM-POP [132] fine-tunes their model using adapters, which are small, trainable networks inserted into each layer of an LLM for interactive planning scenarios.
Alternatively, prompt engineering with in-context learning (ICL) [8] marks a significant advancement in learning from prompts without additional training. Its effectiveness relies on the design and quality of prompts, which can be enhanced with detailed task descriptions, few-shot examples, or model-friendly formats (e.g., ‘###’ as a stop symbol [167]). Moreover, chain-of-thought (CoT) prompting [149] is another emerging approach that incorporates intermediate reasoning steps in prompts. The CoT method substantially enhances the reasoning and problem-solving capabilities of LLMs, making it a dominant technique in robotics applications [163, 86, 128].
@@ 3 Communication
We investigate the utilization of LLMs to facilitate human-like communication in robotics, enabling robots to interact with humans and other robotic agents effectively [97]. We categorize the communication capabilities into two primary areas: (1) language understanding and (2) language generation. We show the detailed categorization in Fig. 1 alongside relevant studies, referred in green cells.
@@ 3.1 Language Understanding
We review language understanding capabilities, addressing how LLMs handle the variability and ambiguity of linguistic inputs through interpretation and grounding processes.
Interpretation transforms natural-language inputs into semantic representations that are easier for robots to process. These representations include formal languages such as linear temporal logic (LTL) [93, 160] and planning domain definition language (PDDL) [155, 89, 18, 42], as well as programming languages such as Python [56, 76]. To aid in interpreting free-form sentences, researchers leverage LLMs’ ICL capabilities, providing guidelines and demonstrations within prompts [122, 56, 89, 76]. Despite the efforts, LLMs often fail to satisfy syntax or capture precise semantics when translating an input into formal languages. To address this issue, researchers suggest simplifying vocabulary or fine-tuning LLMs with domain-agnostic data [93, 160]. For example, Lang2LTL [91] translates landmark-referring expressions in navigational commands into LTL symbols. Further improvements often involve in using human feedback and syntax checkers to correct generated formal language translations [18, 42]. For instance, Guan et al. present a human-in-the-loop translation framework, in which human domain experts repeatedly review PDDL descriptions and provide feedback in natural language [42].
Grounding is another process that maps linguistic expressions to reference targets, such as behaviors or objects, recognizable to robots. Early studies identify mappings that maximize the cosine similarity between word embeddings of LLM outputs and real-world targets [58, 117, 93, 76]. Subsequent studies leverage LLMs’ commonsense knowledge to capture the context of object text labels for improved grounding [118, 41]. For instance, ConceptGraphs [41] demonstrate how LLMs can ground an expression, ‘something to use as a paperweight,’ to a ceramic vase based on size and weight assumptions. However, grounding accuracy depends on the detail and accuracy of the world model. To address this, researchers augment LLMs with multimodal capabilities to directly correlate linguistic inputs with sensory percepts [31, 159, 47, 114], or enable LLMs to interact with environments [158, 168] or humans [61, 120, 109] for better context gathering. For instance, LLM-Grounder [158], a 3D visual grounding method, actively gathers environmental information using vision tools such as LERF [72] and OpenScene [111].
@@ 3.2 Language Generation
Task-dependent language generation focuses on producing language with specific functional objectives, being declarative or imperative. To generate open-ended declarative statements, researchers often provide LLMs with contextual information [20, 96, 62]. However, LLMs often result in repetitive and factually inconsistent outputs, confined by the reliance on previous dialogues and commonsense knowledge [20, 83]. Consequently, researchers augment LLMs with auxiliary knowledge sources to broaden the scope of available information [4, 157, 21]. For instance, Axelsson and Skantze enhance a robot museum guide with knowledge graphs [4]. Furthermore, researchers instruct LLMs to clarify ambiguities by generating imperative instructions requesting human assistance [61, 25]. To improve inference steps, probabilistic models have been introduced to evaluate the uncertainty of situations [120, 109]. For instance, KnowNo [120] and CLARA [109] interaction systems assess confidence and semantic variance, respectively, triggering generation only when these metrics indicate significant uncertainty.
Task-independent language generation involves crafting expressions with social-emotional objectives [11], by embedding non-verbal cues (e.g., non-verbal sounds, hand gestures, and facial expressions) within prompts to enhance engagement and empathy [73, 80]. For example, Khoo et al. have developed a conversational robot that generates empathetic responses using transcribed audio and visual cues [73]. However, conversations with LLMs remain superficial due to the limited knowledge and dialogue history [65]. To overcome this, researchers integrate memory modules into LLMs, enabling them to distill and store information from conversations in a structured format [65, 63, 22, 162]. For example, the companion robot designed by Irfan et al. continuously updates the robot’s memory based on interactions with users to generate personalized dialogues [65].
@@ 4.1 Sensing Modalities
Visual perception tasks involve the interpretation of visual information such as image or point clouds. Pre-trained visual-language models (VLMs), such as CLIP [116] and InstructBLIP [82], allows LLM-based robots to directly utilize image sources. For instance, recent LLM-based manipulation systems, such as TidyBot [152] and RoCo [96], use image-inferred object labels or scene descriptions generated from CLIP and OWL-ViT [100], respectively. In addition, researchers extend reasoning capabilities by applying VLMs on downstream tasks such as image captioning [41] and visual question answering (VQA) [37, 78, 103]. The downstream tasks enable LLMs to subsequently request VLMs to infer object properties (e.g., material, fragility) [37] or ground object parts for grasping [103]. However, images are often challenging to acquire spatial-geometric information.
Alternatively, Huang et al. associate visual-language features from a VLM (i.e., LSeg [81]) with three-dimensional (3D) point clouds for 3D map reconstruction [56]. Further, Jatavallabhula et al. [66] improve this association mechanism with RGB-D images by introducing fine-grained and pixel-aligned features from VLMs. However, association with 3D information tends to be memory intensive, limiting scalability for large scenes [56, 66, 158]. As an alternative solution, researchers often associate geometric and semantic features with 3D scene graphs [41].